# A decomposition of graph $a$-numbers

### v1

@@ Abstract.
In this paper, we establish a combinatorial and topological decomposition formula for the $a$-sequence. As an application, we show that the $a$-sequence is monotone under graph inclusion; that is, $a_{i}(G)\geq a_{i}(H)$ for all $i\geq 0$ whenever $H$ is a subgraph of $G$. We also prove that the $a$-sequence is unimodal in $i$ for a broad class of graphs $G$, including those with a Hamiltonian circuit or a universal vertex. These results provide a new class of topological spaces whose Betti number sequences are unimodal but not necessarily log concave, contributing to the study of real loci in algebraic geometry.
@@
graph $a$-numbers, cohomology, Betti numbers, real toric varieties, graphs, unimodality, log-concavity, blow-ups
@@ 1. Introduction
To present a decomposition formula, we need to introduce some notations. For each pair set $\mathbf{e}$ in $V(G)$, $G+\mathbf{e}$ is the graph obtained from $G$ by adding $\mathbf{e}$ as an edge. If $\mathbf{e}$ is already an edge of $G$, then it does not change the graph, i.e. , $G+\mathbf{e}=G$. We also consider the collection of vertex subsets that induce subgraphs with only even-order connected components, defined as
@@
The second problem deals with the unimodality of the $a_{i}$-numbers as a sequence in $i$. For complex complete toric varieties, it is well known that (even-degree) Betti numbers form a unimodal sequence due to the Hard Lefschetz theorem. [22, 2, 18]. However, this does not generally hold for the real loci of toric varieties. Interestingly, for many graphs $G$, the real toric variety $X^{\mathbb{R}}_{G}$ appears to exhibit unimodal Betti numbers. This phenomenon was posed as an open question in the authors’ previous work [9].
@@
It is also worth noting that these $a$-sequences need not be log-concave as shown in Remark 5.7. While topological spaces with log-concave Betti number sequences have received considerable attention (see, e.g., [17, 3, 15]), much less is known about those that exhibit unimodality without log-concavity. Our results contribute to this direction by exhibiting a broad class of such topological spaces.
The paper is organized as follows. In Section 2, we introduce preliminary concepts. In particular, we introduce the graph associahedra along with their associated complex and real toric varieties, as well as the notion of reconnected complements and certain graph classes that are closed under reconnected complements. Section 3 presents properties of graph $a$-numbers and the motivation behind their decomposition. In Section 4, we prove a decomposition theorem for the $a$-number and, as an application, demonstrate its monotonicity. Section 5 investigates the unimodality of the $a$-sequences. Finally, in Section 6, we study blow-ups of toric varieties, focusing on their real loci. We raise several natural questions and give partial answers concerning how these blow-ups influence the topology of the corresponding real toric varieties.
@@ 2. Reconnected complements
The nestohedron $P_{\mathcal{B}}$ associated with a building set $\mathcal{B}$ naturally lies in $\mathbb{R}^{n+1}$, but is contained in the hyperplane $\sum_{i=1}^{n+1}x_{i}=|\mathcal{B}|$. Modulo the line spanned by the vector $(1,1,\ldots,1)$, this hyperplane can be identified with $\mathbb{R}^{n}$ equipped with the standard lattice $\mathbb{Z}^{n}$. It is well known [19] that the image of any nestohedron under this projection is a Delzant polytope in $\mathbb{R}^{n}$, meaning that at each vertex $v$, the outward normal vectors to the facets containing $v$ can be chosen to form an integral basis of $\mathbb{Z}^{n}$. By the fundamental theorem of toric geometry, $P_{\mathcal{B}}$ induces a smooth projective toric variety $X_{\mathcal{B}}$, called the toric variety associated with $\mathcal{B}$. The fixed point set of $X_{\mathcal{B}}$ under the canonical involution induced by a complex conjugation is denoted by $X^{\mathbb{R}}_{\mathcal{B}}$, and is referred to as the real toric variety associated with $\mathcal{B}$.
@@
Given a finite simple graph $G$, the associated nestohedron $P_{\mathcal{B}(G)}$ is simply denoted by $P_{G}$ and is called a graph associahedron, which is a subclass of nestohedra and has been extensively studied in the literature [6, 25, 12, 8, 14, 7]. The associated complex and real toric varieties, denoted by $X_{G}$ and $X^{\mathbb{R}}_{G}$ respectively, are simply called the complex and real toric varieties associated with $G$.
@@
For each $I\subseteq V(G)$, define the graph $G^{\ast}_{I}$ whose vertex set is $V(G)\setminus I$ and whose edge set consists of unordered pairs $\{a,b\}\subset V(G)\setminus I$ such that there exists a path from $a$ to $b$ in $G_{I\cup\{a,b\}}$. We call $G^{\ast}_{I}$ the reconnected complement of $I$ in $G$.
@@
Let $G$ be the graph obtained from a $6$-cycle graph with vertex set $\{1,2,3,4,5,6\}$ in cyclic order by adding an edge $\{1,4\}$. Let us consider the reconnected complement of $I=\{1,2\}$ in $G$. The edges $\{3,4\}$, $\{4,5\}$, and $\{5,6\}$ remain as edges in $G^{\ast}_{\{1,2\}}$, since they are already present in $G$ and their endpoints are unaffected by the removal. In addition, paths exist from $3$ to $6$ in the induced subgraph $G^{\ast}_{\{1,2,3,6\}}$ and from $4$ to $6$ in $G^{\ast}_{\{1,2,4,6\}}$. In contrast, there is no path from $3$ to $5$ in $G_{\{1,2,3,5\}}$. Hence,
@@ 4. A decomposition formula
$a_{i}(G+\mathbf{e})-a_{i}(G)=\sum_{J\in\mathcal{E}\mathcal{C}(G+\mathbf{e})\setminus\mathcal{E}\mathcal{C}(G)}\left|b(G|_{J})\right|\cdot a_{i-\frac{|J|}{2}}((G+\mathbf{e})^{\ast}_{J}).$
@@ 5. Unimodality of $a$-sequences
We say that a sequence satisfies the tail-peaked condition if it satisfies the following two conditions:
@@
Let $\mathcal{G}$ be a graph class closed under reconnected complement. If the $a$-sequence of each minimal graph in $\mathcal{G}$ satisfies the tail-peaked condition, then the same holds for every graph in $\mathcal{G}$.
@@
Let $G\in\mathcal{G}$ be a simple graph on $m$ vertices. It is immediate that the $a$-sequence of $G$ satisfies the tail-peaked condition when $m\leq 2$. We proceed by induction on $m$. Assume that for some integer $N\geq 2$, the $a$-sequence of every graph in $\mathcal{G}$ with at most $N$ vertices satisfies the tail-peaked condition.
Now let $G\in\mathcal{G}$ be a graph on $N+1$ vertices. If $G$ is minimal in $\mathcal{G}$, then by assumption, its $a$-sequence satisfies the tail-peaked condition. Otherwise, there exists a graph $G^{\prime}\in\mathcal{G}$ with $N+1$ vertices and an edge $\mathbf{e}$ such that $G=G^{\prime}+\mathbf{e}$, where $\mathbf{e}$ is not an edge of $G^{\prime}$. Suppose that the $a$-sequence of $G^{\prime}$ satisfies the tail-peaked condition: that is,
@@
Then, by Theorem 4.5, the $a$-sequence of $G$ also satisfies the tail-peaked condition, completing the inductive step. Therefore, by induction, the $a$-sequence of every graph $G\in\mathcal{G}$ satisfies the tail-peaked condition. ∎
@@
If $G$ is a Hamiltonian graph, then the $a$-sequence satisfies the tail-peaked condition, and hence is unimodal.
@@
Consider the graph class consisting of Hamiltonian graphs together with the single-edge graph. Observe that the cycle graph of order $n\geq 3$ is the unique minimal graph with $n$ vertices in the class. It is known [8] that the $a$-sequence of each cycle graph coincides with the half Pascal triangle, and in particular, it satisfies the tail-peaked condition. Therefore, by Theorem 2.5 (2) and Theorem5.1, we conclude this corollary. ∎
@@
The $a$-sequence of a star graph $G=K_{1,n-1}$ satisfies the tail-peaked condition. In particular,
@@
For each $J\in\mathcal{E}\mathcal{C}(G)\setminus\mathcal{E}\mathcal{C}(G^{\prime})$, we have $v\in J$, so the reconnected complement $G^{\ast}_{J}$ is a complete graph by Theorem 2.5. Hence, by Corollary 5.2, the $a$-sequence of each $G^{\ast}_{J}$ satisfies the tail-peaked condition:
@@
If $G$ has a universal vertex, then the $a$-sequence of $G$ satisfies the tail-peaked condition, and therefore is unimodal.
We turn our attention to the path graph $P_{n}$ on $n$ vertices. It is known [8] that the $a$-sequence of $P_{n}$ corresponds to the rows of the Catalan triangle ([21] A008315). In particular, the sequence satisfies the recurrence
(5.7) $a_{i}(P_{n+1})=a_{i-1}(P_{n})+a_{i}(P_{n}),$
where we set $a_{-1}(P_{n}):=0$. The initial cases can be directly checked: the $a$-sequence of $P_{n}$ is unimodal for small $n$. We claim that the recurrence (5.7) preserves unimodality. Indeed, if the sequence $(a_{0}(P_{n}),a_{1}(P_{n}),\ldots)$ is unimodal, then the entrywise sum with its left-shifted version $(0,a_{0}(P_{n}),a_{1}(P_{n}),\ldots)$ also yields a unimodal sequence. Thus, by induction on $n$, the $a$-sequence of $P_{n}$ is unimodal for all $n\geq 0$. However, the tail-peaked condition does not hold when $n\geq 16$.
We further observe that the $a$-sequence appears to be unimodal for all simple graphs that we have checked. In fact, it was already conjectured in [9] that the $a$-sequence is unimodal for every finite simple graph. It remains an open problem whether this unimodality holds for all finite simple graphs.
@@
Given this apparent unimodality, a natural question is whether the $a$-sequence is also log-concave. However, this does not hold in general. For example, consider the star graph $K_{1,6}$. It is known that the $a$-sequence of $K_{1,6}$ is
$(a_{0}(K_{1,6}),a_{1}(K_{1,6}),\ldots)=(1,6,40,96,\ldots).$
We compute that $(a_{1}(K_{1,6}))^{2}-a_{0}(K_{1,6})a_{2}(K_{1,6})=36-40<0$, which shows that the $a$-sequence of $K_{1,6}$ is not log-concave.
6. Toric blow-ups and real loci

### v2

@@ Abstract.
In this paper, we establish a combinatorial and topological decomposition formula for the $a$-sequence. As an application, we show that the $a$-sequence is monotone under graph inclusion; that is, $a_{i}(G)\geq a_{i}(H)$ for all $i\geq 0$ whenever $H$ is a subgraph of $G$, and obtain the lower and upper bounds of $a_{i}$-numbers. We also prove that the $a$-sequence is unimodal in $i$ for a broad class of graphs $G$, including those with a Hamiltonian circuit or a universal vertex. These results provide a new class of topological spaces whose Betti number sequences are unimodal but not necessarily log concave, contributing to the study of real loci in algebraic geometry.
@@
graph $a$-numbers, cohomology, Betti numbers, real toric varieties, graphs, unimodality, log-concavity, blow-ups, Lefschetz operators
@@ 1. Introduction
When $G$ is a connected graph with an even number of vertices, the $a$-number and $a_{i}$-numbers can also be described via the Möbius function of a poset. Consider the collection of vertex subsets whose induced subgraphs have no odd connected components. This collection, ordered by inclusion, forms a graded poset. The signed $a$-number of $G$ is given by the Möbius function evaluated at the interval from the empty set to the full vertex set. Moreover, the $i$th $a$-number corresponds to the absolute value of the $i$th Whitney number of the first kind of this poset.
@@
To present a decomposition formula, we need to introduce some notations. For each pair set $\mathbf{e}$ in $V(G)$, $G+\mathbf{e}$ is the graph obtained from $G$ by adding $\mathbf{e}$ as an edge. If $\mathbf{e}$ is already an edge of $G$, then it does not change the graph, i.e. , $G+\mathbf{e}=G$. We also consider the collection of vertex subsets that induced subgraphs with only even-order connected components, defined as
@@
As an immediate application of the monotonicity theorem, we obtain the first inequality in the following corollary, by comparing a graph with a spanning tree and the complete graph on the same vertex set. The additional part is more subtle and the proof follows again the decomposition theorem, and we refer to Theorem 5.1.
Corollary 1.3.
Let $G$ be a connected graph on $n$ vertices. Let $K_{n}$ be the complete graph on $n$ vertices, and let $T$ be a spanning tree of $G$. Then, for each $i$,
$a_{i}(T)\leq a_{i}(G)\leq a_{i}(K_{n})=\binom{n}{2i}A_{2i},$
where $A_{k}$ is the $k$th Euler zigzag number. Moreover,
$\frac{n-2i+1}{n-i+1}\binom{n}{i}=a_{i}(P_{n})\leq a_{i}(T)\leq a_{i}(S_{n})=\binom{n-1}{2i-1}A_{2i-1}.$
The second problem deals with the unimodality of the $a_{i}$-numbers as a sequence in $i$. One of the most well-known conjectures related to unimodality is the $g$-conjecture, originally posed by McMullen [24], which characterizes the $h$-vectors of simplicial polytopes. It was resolved by Stanley [30] in the necessity direction, by constructing a projective toric variety whose cohomology ring reflects the $h$-vector, and applying the Hard Lefschetz theorem. This approach laid the foundation for the field of toric topology, which seeks to understand topological invariants via associated combinatorial structures. This unimodality has also been established for more general complex smooth compact toric varieties [2, 23].
However, this does not generally hold for the real loci of toric varieties. In our setting, the $a$-vector of a graph coincides with the rational Betti numbers of a corresponding real toric variety. From the toric topological perspective, it is natural to ask whether these sequences exhibit unimodality, or more broadly, to seek necessary and sufficient conditions for a sequence of positive integers to arise as the $a$-vector of a finite simple graph, as formulated in Problem 6.8. Interestingly, for many graphs $G$, the real toric variety $X^{\mathbb{R}}_{G}$ appears to exhibit unimodal Betti numbers. This phenomenon was posed as an open question in the authors’ previous work [14]. Our decomposition theorem allows us to establish unimodality for several concrete subclasses. We record one such consequence below. See Corollaries 6.2 and 6.6 for the details of the proof.
@@
It is also worth noting that these $a$-sequences need not be log-concave as shown in Remark 6.7. While topological spaces with log-concave Betti number sequences have received considerable attention (see, e.g., [22, 3, 20]), much less is known about those that exhibit unimodality without log-concavity. Although there are [7, 25] several known examples of Betti number sequences of chain complexes that are unimodal but not log-concave, such phenomena have been little studied for Betti number sequences associated with topological spaces. Our results contribute to this direction by exhibiting a broad class of such topological spaces.
The paper is organized as follows. In Section 2, we introduce preliminary concepts. In particular, we introduce the graph associahedra along with their associated complex and real toric varieties, as well as the notion of reconnected complements and certain graph classes that are closed under reconnected complements. Section 3 presents properties of graph $a$-numbers and the motivation behind their decomposition. In Section 4, we prove a decomposition theorem for the $a$-number and, as an application, demonstrate its monotonicity. In Section 5, we prove Corollary 1.3, and Section 6 investigates the unimodality of the $a$-sequences. In Subsection 7.1, we study blow-ups of toric varieties, focusing on their real loci. We raise several natural questions and give partial answers concerning how these blow-ups influence the topology of the corresponding real toric varieties. In Subsection 7.2, we discuss why, in general, the cohomology of real toric varieties does not admit an operator analogous to the Lefschetz operator, even for those arising from graphs.
@@ 2. Reconnected complements
The nestohedron $P_{\mathcal{B}}$ associated with a building set $\mathcal{B}$ naturally lies in $\mathbb{R}^{n+1}$, but is contained in the hyperplane $\sum_{i=1}^{n+1}x_{i}=|\mathcal{B}|$. Modulo the line spanned by the vector $(1,1,\ldots,1)$, this hyperplane can be identified with $\mathbb{R}^{n}$ equipped with the standard lattice $\mathbb{Z}^{n}$. It is well known [27] that the image of any nestohedron under this projection is a Delzant polytope in $\mathbb{R}^{n}$, meaning that at each vertex $v$, the outward normal vectors to the facets containing $v$ can be chosen to form an integral basis of $\mathbb{Z}^{n}$. By the fundamental theorem of toric geometry, $P_{\mathcal{B}}$ induces a smooth projective toric variety $X_{P_{\mathcal{B}}}$, called the toric variety associated with $\mathcal{B}$. The fixed point set of $X_{P_{\mathcal{B}}}$ under the canonical involution induced by a complex conjugation is denoted by $X^{\mathbb{R}}_{P_{\mathcal{B}}}$, and is referred to as the real toric variety associated with $\mathcal{B}$.
@@
Given a finite simple graph $G$, the associated nestohedron $P_{P_{\mathcal{B}}(G)}$ is simply denoted by $P_{G}$ and is called a graph associahedron, which is a subclass of nestohedra and has been extensively studied in the literature [8, 33, 26, 17, 10, 19, 9]. The associated complex and real toric varieties, denoted by $X_{G}$ and $X^{\mathbb{R}}_{G}$ respectively, are simply called the complex and real toric varieties associated with $G$.
@@
For each $I\subseteq V(G)$, define the graph $G^{\ast}_{I}$ whose vertex set is $V(G)\setminus I$ and whose edge set consists of unordered pairs $\{a,b\}\subset V(G)\setminus I$ such that there exists a path from $a$ to $b$ in $G|_{I\cup\{a,b\}}$. We call $G^{\ast}_{I}$ the reconnected complement of $I$ in $G$.
@@
Let $G$ be the graph obtained from a $6$-cycle graph with vertex set $\{1,2,3,4,5,6\}$ in cyclic order by adding an edge $\{1,4\}$. Let us consider the reconnected complement of $I=\{1,2\}$ in $G$. The edges $\{3,4\}$, $\{4,5\}$, and $\{5,6\}$ remain as edges in $G^{\ast}_{\{1,2\}}$, since they are already present in $G$ and their endpoints are unaffected by the removal. In addition, paths exist from $3$ to $6$ in the induced subgraph $G^{\ast}_{\{1,2,3,6\}}$ and from $4$ to $6$ in $G^{\ast}_{\{1,2,4,6\}}$. In contrast, there is no path from $3$ to $5$ in $G|_{\{1,2,3,5\}}$. Hence,
@@ 4. A decomposition formula
(4.2) $a_{i}(G+\mathbf{e})-a_{i}(G)=\sum_{J\in\mathcal{E}\mathcal{C}(G+\mathbf{e})\setminus\mathcal{E}\mathcal{C}(G)}\left|b(G|_{J})\right|\cdot a_{i-\frac{|J|}{2}}((G+\mathbf{e})^{\ast}_{J}).$
@@
Example 4.6.
Let
$C(n,k)=\binom{n+k}{k}-\binom{n+k}{k-1}=\frac{n-k+1}{n+1}\binom{n+k}{k}$.
It is well known that $C(n,k)$ counts the strings with $n$ letters $X$ and $k$ letters $Y$ such that no initial segment of the string contains more $Y$’s than $X$’s ([29] A008315). These numbers form the entries of the Catalan triangle, and $C(n,n)$ is the $n$th Catalan number, denoted by $\mathcal{C}_{n}$. For the path graph $P_{n}$, it was shown in [10] that, for $0\leq i\leq\lfloor(n+1)/2\rfloor$,
$a_{i}(P_{n})=C(n-i,i).$
In this example, we reprove this formula using the decomposition formula (4.2).
We regard $P_{n}$ as the graph on $n+1]=\{1,\ldots,n+1\}$ with edges $\{i,i+1\}$ for $i=1,\ldots,n-1$, so that its $a_{i}$-numbers agree with those of the usual path graph on $n$ vertices. Let $e=\{n,n+1\}$ and observe that $P_{n}+e=P_{n+1}$. We now describe the subsets $J$ appearing as non-zero terms on the right-hand side in (4.2). By Proposition 3.3 (2), the admissible subsets $J$ are exactly those subsets of $[n+1]$ that contain $n+1$ and consist of consecutive vertices of even cardinality. Writing $|J|=2j$, such a subset is of the form
$J=\{n+2-2j,\ldots,n+1\}.$
In this case, one easily checks that
$P_{n}|_{J}\cong P_{2j-1}\quad\text{ and }\quad(P_{n+1})^{\ast}_{J}\cong P_{n+1-2j}.$
We proceed by induction on $n$. Assume that for all $m\leq n$ and all $i$,
$a_{i}(P_{m})=C(m-i,i).$
Then, for $2j-1\leq n$, we have
$b(P_{2j-1})=\mathcal{C}_{j-1}.$
Using the induction hypothesis, the decomposition formula yields
$a_{i}(P_{n+1})$ $=C(n-i,i)+\sum_{j=1}^{i}\mathcal{C}_{j-1}C(n+1-i-j,i-j)$
$=C(n-i,i)+\sum_{j=0}^{i}\mathcal{C}_{j}C(n-i-j,i-j-1)$
$=C(n-i,i)+C(n+1-i,i-1)$
$=C(n+1-i,i),$
which completes the proof.
@@
5. Upper and Lower bounds of $a_{i}$-numbers
In the previous section, we discussed the monotonicity of the $a_{i}$-numbers. In this section, we use this property to derive bounds for the $a_{i}$-numbers and thereby prove Corollary 1.3. By Corollary 4.7, the $a_{i}$-number of a graph $G$ is sandwiched between the $a_{i}$-number of a spanning tree of $G$ and that of the complete graph on the same number of vertices. Therefore, to complete the proof of Corollary 1.3, it suffices to show that for an arbitrary tree $T$, its $a_{i}$-number lies between those of the path graph and the star graph with the same number of vertices.
Throughout this section, the path graph and the star graph on $n$ vertices are denoted by $P_{n}$ and $K_{1,n-1}$, respectively. It is known that for $0\leq i\leq\lfloor(n+1)/2\rfloor$,
$a_{i}(P_{n})=\frac{n-2i+1}{n-i+1}\binom{n}{i},\quad\text{ and }\quad a_{i}(K_{1,n-1})=\binom{n-1}{2i-1}A_{2i-1},$
where $A_{k}$ denotes the $k$th Euler zigzag number. See Example 4.6 and [10] for details.
Theorem 5.1.
Let $T_{n}$ be a tree on $n$ vertices. Then, for each non-negative integer $i$,
(5.1) $a_{i}(P_{n})\leq a_{i}(T_{n})\leq a_{i}(K_{1,n-1}).$
To prove the above theorem, we prepare the following lemma.
Lemma 5.2.
For any tree $T_{n}$ on $n$ vertices,
$|b(P_{n})|\leq|b(T_{n})|\leq|b(K_{1,n-1})|.$
Proof.
Let $G$ be a graph on $n$ vertices, and let $X_{G}^{\mathbb{R}}$ be the real toric variety associated with $G$. By definition, $b(G)$ is the alternating sum of the $a_{i}$-numbers, and $a_{i}(G)=\beta_{i}(X_{G}^{\mathbb{R}};\mathbb{Q})$, where
$\beta_{i}(X;\mathbb{F})=\dim_{\mathbb{F}}H_{i}(X;\mathbb{F})$
denotes the $i$th Betti number of a space $X$ with coefficients in a field $\mathbb{F}$. Thus
$b(G)$ $=\sum_{i\geq 0}(-1)^{i}\beta_{i}(X_{G}^{\mathbb{R}};\mathbb{Q})=\chi(X_{G}^{\mathbb{R}})$
$=\sum_{j\geq 0}(-1)^{j}\beta_{j}(X_{G}^{\mathbb{R}};\mathbb{Z}_{2}),$
where $\chi(X)$ denotes the Euler characteristic of a space $X$.
Assume that $n=2k+1$. Let $(h_{0},\ldots,h_{2k})$ be the $h$-vector of the graph associahedron $P_{G}$, and let $\gamma=(\gamma_{0},\ldots,\gamma_{k})$ be the gamma vector of $P_{G}$ defined by
$\sum_{i=0}^{2k}h_{i}t^{i}=\sum_{i=0}^{k}\gamma_{i}t^{i}(1+t)^{2k-2i}.$
By [16], we have $\beta_{j}(X_{G}^{\mathbb{R}};\mathbb{Z}_{2})=h_{j}$ for all $j$, hence
$b(G)=\sum_{j=0}^{2k}(-1)^{j}h_{j}.$
Evaluating the defining identity at t=1t=-1 gives $\sum_{j=0}^{2k}(-1)^{j}h_{j}=(-1)^{k}\gamma_{k}$, hence
$|b(G)|=\gamma_{k}.$
By [6], for every $i$,
$\gamma_{i}(P_{P_{n}})\leq\gamma_{i}(P_{T_{n}})\leq\gamma_{i}(P_{K_{1,n-1}}).$
Applying this with $i=k$ and using $|b(G)|=\gamma_{k}$ proves the claim for $n=2k+1$. If $n$ is even, then $b(G)=0$ by Proposition 3.3, so the inequality holds. ∎
Proof of Theorem 5.1.
We will prove the theorem by mathematical induction on $n$. It is straightforward to verify that (5.1) holds for all $n\leq 4$. Assume that (5.1) holds for all $n\leq N$. A tree $T_{N+1}$ has at least two leaves. Let $v$ be one such leaf, and let $u$ denote the vertex adjacent to $v$ in $T_{N+1}$. Define $F_{N}$ to be the forest obtained from $T_{N+1}$ by deleting the edge $e=(u,v)$, i.e.
$T_{N+1}=F_{N}+e.$
Note that
$a_{i}(P_{N})\leq a_{i}(F_{N})\leq a_{i}(K_{1,N-1}).$
Applying Theorem 4.5, we have
(5.2) $a_{i}(T_{N+1})-a_{i}(F_{N})=\sum_{J\in\mathcal{E}\mathcal{C}(T_{N+1})\setminus\mathcal{E}\mathcal{C}(F_{N})}\left|b(F_{N}|_{J})\right|\cdot a_{i-\frac{|J|}{2}}((T_{N+1})^{\ast}_{J}).$
The difference $\mathcal{E}\mathcal{C}(T_{N+1})\setminus\mathcal{E}\mathcal{C}(F_{N})$ consists precisely of those sets of the form $J=J^{\prime}\cup\{v\}$, where $F_{N}|_{J^{\prime}}$ is a connected subgraph of odd order. For any $J\in\mathcal{E}\mathcal{C}(T_{N+1})\setminus\mathcal{E}\mathcal{C}(F_{N})$, we have the following observations.
  1. (1)
    $b(F_{N}|_{J})=b(P_{|J|})$ and $(T_{N+1})^{\ast}_{J}=P_{N+1-J}$ when $T_{N+1}$ is a path graph.
  2. (2)
    $b(F_{N}|_{J})=b(K_{|J|-1,1})$ and $(T_{N+1})^{\ast}_{J}=K_{N+1-J}$, when $T_{N+1}$ is a star graph.
Since $T_{N+1}$ is connected, for each $k\geq 0$, there is at least one such $J^{\prime}$ with $|J^{\prime}|=2k+1$. For such subsets $J$ appearing in (5.2), the number of admissible choices is minimized when $T_{N+1}$ is a path graph and maximized when $T_{N+1}$ is a star graph for each even cardinality of $J$. Hence, the lower and upper bounds of the right-hand side of (5.2) are attained when $T_{N+1}$ is a path graph and a star graph, respectively. Combining this with Lemma 5.2 and the induction hypothesis, we complete the proof of the theorem. ∎
@@ 6. Unimodality of $a$-sequences
We say that a sequence is penultimately increasing if it satisfies the following two conditions:
@@
Let $\mathcal{G}$ be a graph class closed under reconnected complement. If the $a$-sequence of each minimal graph in $\mathcal{G}$ is penultimately increasing, then the same holds for every graph in $\mathcal{G}$.
@@
Let $G\in\mathcal{G}$ be a simple graph on $m$ vertices. It is immediate that the $a$-sequence of $G$ is penultimately increasing when $m\leq 2$. We proceed by induction on $m$. Assume that for some integer $N\geq 2$, the $a$-sequence of every graph in $\mathcal{G}$ with at most $N$ vertices is penultimately increasing.
Now let $G\in\mathcal{G}$ be a graph on $N+1$ vertices. If $G$ is minimal in $\mathcal{G}$, then by assumption, its $a$-sequence is penultimately increasing. Otherwise, there exists a graph $G^{\prime}\in\mathcal{G}$ with $N+1$ vertices and an edge $\mathbf{e}$ such that $G=G^{\prime}+\mathbf{e}$, where $\mathbf{e}$ is not an edge of $G^{\prime}$. Suppose that the $a$-sequence of $G^{\prime}$ is penultimately increasing: that is,
@@
Then, by Theorem 4.5, the $a$-sequence of $G$ is also penultimately increasing, completing the inductive step. Therefore, by induction, the $a$-sequence of every graph $G\in\mathcal{G}$ is penultimately increasing. ∎
@@
If $G$ is a Hamiltonian graph, then the $a$-sequence is penultimately increasing, and hence is unimodal.
@@
Consider the graph class consisting of Hamiltonian graphs together with the single-edge graph. Observe that the cycle graph of order $n\geq 3$ is the unique minimal graph with $n$ vertices in the class. It is known [10] that the $a$-sequence of each cycle graph coincides with the half Pascal triangle, and in particular, it is penultimately increasing. Therefore, by Theorem 2.5 (2) and Theorem6.1, we conclude this corollary. ∎
@@
The $a$-sequence of a star graph $G=K_{1,n-1}$ is penultimately increasing. In particular,
@@
For each $J\in\mathcal{E}\mathcal{C}(G)\setminus\mathcal{E}\mathcal{C}(G^{\prime})$, we have $v\in J$, so the reconnected complement $G^{\ast}_{J}$ is a complete graph by Theorem 2.5. Hence, by Corollary 6.2, the $a$-sequence of each $G^{\ast}_{J}$ is penultimately increasing:
@@
If $G$ has a universal vertex, then the $a$-sequence of $G$ is penultimately increasing, and therefore is unimodal.
It is important to note that not all graphs have a penultimately increasing $a$-vector. We turn our attention to the path graph $P_{n}$ on $n$ vertices. As in Example 4.6, we have $a_{i}(P_{n})=\frac{n-2i+1}{n-i+1}\binom{n}{i}$. Hence the consecutive ratio is
$r_{i}:=\frac{a_{i+1}}{a_{i}}=\frac{(n-2i-1)(n-i+1)}{(i+1)(n-2i+1)}.$
A direct computation shows that $r_{i}$ is strictly decreasing in $i$ on the range $0\leq i<\lfloor s/2\rfloor$ (equivalently, $r_{i}-r_{i+1}>0$ for all such $i$). Therefore the sequence $a_{i}(P_{n})$ is log-concave for each $n$, and in particular unimodal. Moreover, since $r_{i}\geq 1$ if and only if
$i\leq\frac{n-\sqrt{n+2}}{2},$
the sequence increases up to an index $i$ close to $(n-\sqrt{n+2})/2$ and then decreases. Hence, it is not penultimately increasing for sufficiently large $n$. In fact, this property fails for $n\geq 16$.
To the best of our knowledge, there is no known example of a graph whose $a$-vector fails to be unimodal. In fact, it was already conjectured in [14] that the $a$-sequence is unimodal for every finite simple graph.
@@
Given this apparent unimodality, a natural question is whether the $a$-sequence is also log-concave. However, this does not hold in general. For example, consider the star graph $K_{1,6}$, whose $a$-sequence is given by
$(a_{0},a_{1},a_{2},a_{3},a_{4},\ldots)=(1,6,40,96,0,\ldots).$
This sequence is penultimately increasing and hence unimodal, but it fails to be log-concave, as $a_{1}^{2}-a_{0}a_{2}=-4<0$.
Problem 6.8.
Characterize the necessary and sufficient conditions for a sequence of positive integers to be realized as the $a$-sequence of a finite simple graph.
7. Further geometric remarks
7.1. Toric blow-ups and real loci